# Aperiodic points for outer billiards

### v3

@@ 5.1. Boundedness
The proof of Theorem 5.2 now becomes a straightforward consequence of Baire’s theorem (see, e.g. Section 7.3 of [KF75]): if $Y$ is a compact metric space, then any countable union of closed nowhere dense subsets of $Y$ is nowhere dense, in particular, is different from $Y$.
@@
It suffices to show that $Y^{\xi}_{\alpha}\subset X^{\xi}_{\alpha}$ by transfinite induction on $\alpha$. The induction base is given as an assumption. Next, if $Y^{\xi}_{\alpha}\subset X^{\xi}_{\alpha}$, then any point of $Y^{\xi}_{\alpha}$, where the restriction of $\xi$ to $X^{\xi}_{\alpha}$ is locally unbounded, is also a point of $X^{\xi}_{\alpha}$, at which the restriction of $\xi$ to $X^{\xi}_{\alpha}$ is locally unbounded. Hence, $Y^{\xi}_{\alpha+1}\subset X^{\xi}_{\alpha+1}$. Passing to the intersections also preserves the inclusion, and the lemma now follows by transfinite induction. ∎
@@ 5.2. Centers
The set $X_{0}$ is closed by definition, and the intersection of closed sets is closed. It is therefore enough to show that $X_{\alpha+1}$ is closed assuming that so is $X_{\alpha}$. Consider a sequence $x_{n}\in X_{\alpha+1}$ converging to some point $x\in X$. We want to prove that $x\in X_{\alpha+1}$. Pick a neighborhood $O_{x}$ of $x$; then all $x_{n}$ with sufficiently large $n$ lie in $O_{x}$. By definition of $X_{\alpha+1}$ and since $O_{x}$ is also a neighborhood of $x_{n}\in X_{\alpha+1}$ with large $n$, there is a point $y\in O_{x}\cap X_{\alpha}$ and a positive integer $m$ such that $h^{m}(y)\in O_{x}$. The existence of such $y$ proves that $x\in X_{\alpha+1}$, as claimed. We obtain that all $X_{\alpha}$ are closed, hence compact.
@@
If ordinals $\alpha$ are sufficiently large, then all $X_{\alpha}$ are the same; denote them by $Y$. By Lemma 5.7, the set $Y$ is compact and nonempty. Also, by definition of $X_{\alpha}$ and since $X_{\alpha+1}=X_{\alpha}$, the set $Y$ satisfies the following property: for every $x\in Y$ and any neighborhood $O_{x}$ of $x$, there is a point $y\in O_{x}$ and a positive integer $n$ with $h^{n}(y)\in O_{x}$. Either $x$ is not isolated, or $y=x=h^{n}(y)$ provided that $O_{x}$ is sufficiently tight. The latter case, $y$ is periodic, a contradiction with our assumption. ∎
@@ 6. Existence of aperiodic points
Here, the assumption that $f$ is unbranched can probably be removed; however, it simplifies the arguments, and it does hold for polygon exchange transformations arising from outer billiards on regular polygons. Whether the assumption about a dynamic Hadwiger invariant being nonzero is essential, we do not know. Theorem 6.1, together with Corollary 3.7, imply Theorem 1.1, since the map $g_{X}$ from Section 3 is clearly unbranched and has non-vanishing dynamic Hadwiger invariants. Our standing assumption through the proof of Theorem 6.1 will be that $f$ is a non-periodic polygon exchange transformation, however, it has no aperiodic points.
@@ 6.1. Bounded levels
Recall that $\mathcal{J}$ is the ideal Julia set of $f$, and that $h_{f}:\mathcal{J}\to\mathcal{J}$ is the homeomorphism induced by $f$; it acts on $\mathcal{J}$ without periodic points. Since there are no aperiodic points of $f$, the bilevel function $\mathrm{bilev}$ is everywhere finite on $J$. Define the corresponding function $\xi:\mathcal{J}\to\mathbb{Z}_{\geqslant 0}$ as follows: the value of $\xi$ at an essential nest $(Q_{n})$ equals the maximum value of the function $\mathrm{bilev}$ on the impression of this nest. Observe that either the impression of $(Q_{n})$ is a single point, in which case there is only one value of $\mathrm{bilev}$, or the impression lies in $\Gamma_{k}\cup\Gamma_{-k}$ for some $k>0$, in which case the function $\mathrm{bilev}$ is bounded above by $k$. Thus, in any case $\xi$ is well defined on $\mathcal{J}$. It is also easy to see that $\xi$ is lower-semicontinious: this follows from semi-continuity of $\mathrm{bilev}$, see Lemma 2.10.
@@
We do not claim that $J^{L_{1}}_{*}$ and $J^{L_{2}}_{*}$ are disjoint when $L_{1}$ and $L_{2}$ are transverse.
@@
Any point $z\in J^{L}_{*}$ lies in the impression of some nest $(Q_{n})$ supported by a line through $z$ parallel to $L$. The image $(Q^{\prime}_{n})$ of this nest under $h_{f}$ contains the point $f_{L}(z)$ provided that $f_{L}$ is defined on $z$. It follows that $f_{L}(z)\in J^{L}_{*}$. Note, finally, that $f_{L}$ can be undefined at $z$ only if $z$ is one of the finitely many vertices of $\Gamma_{1,L}$. Finally, the fact that points of $J^{L}_{*}$ are nonwandering under the iterates of $f_{L}$ follows from the corresponding property of $h_{f}$ and Lemma 2.15. ∎
@@ 6.2. Geometry and dynamics of $J^{L}_{*}$
In this section, we discuss what the set $J^{L}_{*}$ may look like and how $f_{L}$ may act on this set. Assume that $J^{L}_{*}\neq\varnothing$ for a given choice of a co-oriented line $L$. We will think of $L$ as being a horizontal line, and will refer to the positive side of $L$ as the upper side. Consider the union $I_{L}$ of all edges of $\Gamma_{1,L}$ that contain points of $J^{L}_{*}$ inside, i.e., not only as endpoints. Then $f_{L}$ acts on $I_{L}$ as a translation scissor congruence. It can be extended to an interval exchange transformation $\varphi_{L}$ of $\Gamma_{1,L}$. Note that the dynamics of $\varphi_{L}$ and $f_{L}$ must match on $J^{L}_{*}$ but may be different otherwise. We now need some general properties of interval exchange maps.
@@
Let $\varepsilon>0$ be a sufficiently small number, in particular, smaller than the distance between any two distinct lines containing components of $J^{L}_{*}$. A forward trapezoid is defined as the trapezoid with horizontal bases such that the bottom base is a forward 1-cell of $J^{L}_{*}$, the height is $\varepsilon$, and, near the endpoints of its bottom base, the trapezoid coincides with elementary sectors. When $\varepsilon$ needs to be indicated, we speak of forward $\varepsilon$-trapezoids. Backward trapezoids are defined in the same way, replacing forward 1-cells with backward 1-cells. The union of the forward trapezoids is called the forward $\varepsilon$-train of $J_{L*}$, or just the forward train of $J_{L*}$ if $\varepsilon$ is fixed. Backward trains are defined similarly.
@@
Note that the $\varepsilon$-trains of $J^{L}_{*}$, both forward and backward, are contained in an arbitrarily small neighborhood of $J^{L}_{*}$, provided that $\varepsilon$ is small enough for the chosen neighborhood. The set theoretic difference between the backward train and the forward train, with the same $\varepsilon$, consists of several triangles (possibly degenerate) cut out of the union of the two trains by $\Gamma_{1}$. Similarly, the forward train minus the backward train is a union of finitely many triangles cut out of the union of the two trains by $\Gamma_{-1}$.
@@
If $f$ is unbranched, then there can be no non-degenerate triangles cut out by $\Gamma_{1}$ or by $\Gamma_{-1}$. Therefore, the forward and the backward trains are subdivision of the same polygon, and $f$ acts on this polygon as a polygon exchange transformation. Moreover, the restriction of this polygon exchange transformation to every component of its domain is a parallel translation by a vertor parallel to $L$, that is, by a horizontal vector.
@@
Suppose that an essential nest $(Q_{n})$ belongs to $\mathcal{J}_{*}$, and fix $\varepsilon>0$. There is a co-oriented line $L$ containing an edge of $\Gamma_{1}$ and such that, for all sufficiently large $n$, the bi-piece $Q_{n}$ is contained in the $\varepsilon$-train of $J^{L}_{*}$.
@@ 6.3. Horizontal polygon exchange transformations
Fix a co-oriented line $L$ and, as before, refer to its direction as the horizontal direction, and to its positive side as the upper side. Suppose that a polygon exchange transformation $f_{\Theta}$ of a polygon $\Theta$ translates all level 1 pieces by vectors parallel to $L$. In this case, $f_{\Theta}$ is called a horizontal polygon exchange transformation with respect to $L$. Our standing assumption in Section 6.3 is that $f_{\Theta}$ is a horizontal polygon exchange transformation with respect to $L$. There are several special properties of $f_{\Theta}$. Any horizontal line $L_{t}$ parallel to $L$ intersects $\Theta$ over a finite union of horizontal intervals. Note that $f_{\Theta}$ induces an interval exchange transformation of $\Theta\cap L_{t}$. Here, the parameter $t$ can be understood e.g. as a height over $L=L_{0}$, and we write $\varphi_{t}$ as the interval exchange transformation of $\Theta\cap L_{t}$ induced by $f_{\Theta}$.
@@
Before giving a proof of Theorem 6.13, we make several preparatory remarks. There are finitely many lines $L_{t_{0}}$, $\dots$, $L_{t_{m}}$ such that the vertices of all level 1 pieces of $f_{\Theta}$ are contained in the union of these lines. Assume that the set of lines $L_{t_{i}}$ with the just mentioned property is chosen to be minimal by inclusion. In this case, the part $\Theta_{i}$ of $\Theta$ between $L_{t_{i}}$ and $L_{t_{i+1}}$ is a union of trapezoids with horizontal edges in $L_{t_{i}}$ and $L_{t_{i+1}}$. Some of these edges may degenerate to points, letting the corresponding trepezoids degenerate to triangles. Call the thus obtained trapezoids between $L_{t_{i}}$ and $L_{t_{i+1}}$ the trapezoids on the $i$-th floor. Observe that $f_{\Theta}$ acts as a horizontal polygon exchange transformation on $\Theta_{i}$. Knowing that the dynamical invariant $\mathrm{Inv}_{L}$ is additive with respect to polygonal subdivisions of the domain, it suffices to prove Theorem 6.13 for the union of all trapezoids on a given floor. Moreover, if we consider the first return map to any given trapezoid and then subdivide it further into floors if necessary, then we reduce Theorem 6.13 to the case when $\Theta$ is just an $L$-trapezoid, i.e., a trapezoid with top and bottom edges, possibly degenerate, parallel to $L$. We therefore assume now that $\Theta$ is an $L$-trapezoid.
@@
where $x\wedge y$ means $x\otimes y-y\otimes x$, and all tensor products are over $\mathbb{Q}$.
@@
Write $\lambda_{i}(t)$ as $\lambda_{i}(0)+t\lambda^{\prime}_{i}(0)$ and substitute this expression into the formula $\sum\lambda_{i}(t)\otimes\tau_{i}$ for the invariant $\mathrm{Inv}(\varphi_{t})$. Recall that $\mathbb{R}\otimes\mathbb{R}$ can be viewed as a vector space over $\mathbb{R}$, with left multiplication by real scalars satisfying the identity $(ta)\otimes b=t(a\times b)$. It follows that $\mathrm{Inv}(\varphi_{t})$ can be written as the sum of two terms: the first term is independent of $t$ and is equal to $\mathrm{Inv}(\varphi_{0})$, whereas the second term is $t$ times the sum of $\lambda^{\prime}_{i}(0)\otimes\tau_{i}$ over $i$ ranging from 1 to $m$. Clearly, it suffices to show that this second term, equal to $\frac{d}{dt}\mathrm{Inv}(\varphi_{t})|_{t=0}$, vanishes. Substituting the expression for $\tau$ as $\tau=\Omega\lambda(0)$, we obtain that
@@
Here, the factor $\Omega_{ij}$ has been moved to the front of the tensor product since $\Omega_{ij}\in\mathbb{Q}$. The left-hand side is a time derivative of a function with values in a vector space. As the function is linear, the derivative is well defined even without introducing any norm.
@@
By the above, it suffices to assume that $\Theta$ is an $L$-trapezoid between $L_{0}$ and $L_{1}$, so that $\varphi_{t}$ for $t\in[0;1]$ form a linear one parameter family of interval exchange transformations. Since all translation vectors are horizontal, the space $\mathbb{R}^{2}$ of translations reduces to $\mathbb{R}^{1}$, where all $\tau_{i}(t)$ lie. With this identification, the dynamical Hadwiger invariant of $f_{\Theta}$ identifies with the difference $\mathrm{Inv}(\varphi_{0})-\mathrm{Inv}(\varphi_{1})$ of two SAF invariants. By Lemma 6.15, the latter difference is zero. ∎
As it turns out, other dynamical Hadwiger invariants of $f_{\Theta}$ vanish as well.
@@
All dynamical Hadwiger invariants of a horizontal polygon exchange transformation $f_{\Theta}$ are zero.
@@
Vanishing of $\mathrm{Inv}_{L}(f_{\Theta})$ follows from Theorem 6.13. Let $L^{\prime}$ be a line transverse to $L$; we want to show that $\mathrm{Inv}_{L^{\prime}}(f_{\Theta})=0$. It suffices to assume that $\Theta$ is an $L$-trapezoid between $L_{0}$ and $L_{1}$. Also, one may assume without loss of generality that the positive side of $L^{\prime}$ is on the right of $L^{\prime}$. Write $\ell^{\prime}$ for the length of the intersection of the strip between $L_{0}$ and $L_{1}$ with any line parallel to $L^{\prime}$. Suppose that $i_{1}<\dots<i_{k}$ are all the values of the index $i$, for which the left end of $I_{i}(t)$ for $10<t<$ 1 lies on an edge of $\Gamma_{1}(f_{\Theta})$ parallel to $L^{\prime}$. One also writes $j_{1}<\dots<j_{l}$ for the values of $j$ such that the right end of $I_{j}(t)$ for $10<t<$ 1 lies on an edge of $\Gamma_{1}(f_{\Theta})$ parallel to $L^{\prime}$. The value $\mathrm{Inv}_{L^{\prime}}(f_{\Theta})$ is by definition equal to
@@
All the dynamical Hadwiger invariants of $f_{\circ}$ are the same as those of $f$. Hence, at least one of these invariants takes a nonzero value on $f_{\circ}$.
@@
Indeed, this follows from the additivity of the invariants and the fact that the restriction of $f$ on every train has all the dynamical Hadwiger invariants zero, by Theorem 6.16. ∎
Since $f_{\circ}$ has a nonzero dynamical Hadwiger invariant, it is not periodic. However, $f_{\circ}$ cannot have aperiodic points — all aperiodic points of $f_{\circ}$ would also be aperiodic points of $f$. The ideal Julia set $\mathcal{J}(f_{\circ})$ of $f_{\circ}$ and the essential part of it, $\mathcal{J}_{*}(f_{\circ})$, are therefore nonempty.

### v4

@@ 5.1. Boundedness
The proof of Theorem 5.2 now becomes a straightforward consequence of Baire’s theorem (see, e.g. Section 7.3 of [KF75]): if $Y$ is a complete metric space, then any countable union of closed nowhere dense subsets of $Y$ has empty interior, in particular, cannot be $Y$.
@@
It suffices to show that $Y^{\xi}_{\alpha}\subset X^{\xi}_{\alpha}$ by transfinite induction on $\alpha$. The induction base is given as an assumption. Next, if $Y^{\xi}_{\alpha}\subset X^{\xi}_{\alpha}$, then any point of $Y^{\xi}_{\alpha}$, where the restriction of $\xi$ to $Y^{\xi}_{\alpha}$ is locally unbounded, is also a point of $X^{\xi}_{\alpha}$, at which the restriction of $\xi$ to $X^{\xi}_{\alpha}$ is locally unbounded. Hence, $Y^{\xi}_{\alpha+1}\subset X^{\xi}_{\alpha+1}$. Passing to the intersections also preserves the inclusion, and the lemma now follows by transfinite induction. ∎
@@ 5.2. Centers
The set $X_{0}$ is closed by definition, and the intersection of closed sets is closed. It is therefore enough to show that $X_{\alpha+1}$ is closed assuming that so is $X_{\alpha}$. Consider a sequence $x_{n}\in X_{\alpha+1}$ converging to some point $x\in X$. We want to prove that $x\in X_{\alpha+1}$. Pick a neighborhood $O_{x}$ of $x$ in $X_{\alpha}$; then all $x_{n}$ with sufficiently large $n$ lie in $O_{x}$. By definition of $X_{\alpha+1}$ and since $O_{x}$ is also a neighborhood of $x_{n}\in X_{\alpha+1}$ with large $n$, there is a point $y\in O_{x}\cap X_{\alpha}$ and a positive integer $m$ such that $h^{m}(y)\in O_{x}$. The existence of such $y$ proves that $x\in X_{\alpha+1}$, as claimed. We obtain that all $X_{\alpha}$ are closed, hence compact.
@@
If ordinals $\alpha$ are sufficiently large, then all $X_{\alpha}$ are the same; denote them by $Y$. By Lemma 5.7, the set $Y$ is compact and nonempty. Also, by definition of $X_{\alpha}$ and since $X_{\alpha+1}=X_{\alpha}$, the set $Y$ satisfies the following property: for every $x\in Y$ and any neighborhood $O_{x}$ of $x$, there is a point $y\in O_{x}$ and a positive integer $n$ with $h^{n}(y)\in O_{x}$. Either $x$ is not isolated, or $y=x=h^{n}(y)$ provided that $O_{x}$ is sufficiently tight. The latter case means $y$ is periodic, which is a contradiction with our assumption. ∎
@@ 6. Existence of aperiodic points
Here, the assumption that $f$ is unbranched can probably be removed; however, it simplifies the arguments, and it does hold for polygon exchange transformations arising from outer billiards on regular polygons. Whether the assumption about a dynamic Hadwiger invariant being nonzero is essential, we do not know. Theorem 6.1, together with Corollary 3.7, imply Theorem 1.1, since the map $g_{X}$ from Section 3 is clearly unbranched and has non-vanishing dynamic Hadwiger invariants. Our standing assumption throughout the proof of Theorem 6.1 will be that $f$ is a non-periodic polygon exchange transformation, however, it has no aperiodic points.
@@ 6.1. Bounded levels
Recall that $\mathcal{J}$ is the ideal Julia set of $f$, and that $h_{f}:\mathcal{J}\to\mathcal{J}$ is the homeomorphism induced by $f$; it acts on $\mathcal{J}$ without periodic points. Since there are no aperiodic points of $f$, the bilevel function $\mathrm{bilev}$ is everywhere finite on $J$. Define the corresponding function $\xi:\mathcal{J}\to\mathbb{Z}_{\geqslant 0}$ as follows: the value of $\xi$ at an essential nest $(Q_{n})$ equals the maximum value of the function $\mathrm{bilev}$ on the impression of this nest. Observe that either the impression of $(Q_{n})$ is a single point, in which case there is only one value of $\mathrm{bilev}$, or the impression lies in $\Gamma_{k}\cup\Gamma_{-k}$ for some $k>0$, in which case the function $\mathrm{bilev}$ is bounded above by $k$. Here, we used our standing assumption that $f$ has no aperiodic points. Thus, in any case $\xi$ is well defined on $\mathcal{J}$. It is also easy to see that $\xi$ is lower-semicontinious: this follows from semi-continuity of $\mathrm{bilev}$, see Lemma 2.10.
@@
We do not claim that $J^{L_{1}}_{*}$ and $J^{L_{2}}_{*}$ are disjoint when $L_{1}$ and $L_{2}$ are transverse; the claim of the lemma concerns only the sets $\mathcal{J}^{L_{1}}_{*}$ and $\mathcal{J}^{L_{2}}_{*}$.
@@
Any point $z\in J^{L}_{*}$ lies in the impression of some nest $(Q_{n})$ supported by a line through $z$ parallel to $L$. The image $(Q^{\prime}_{n})$ of this nest under $h_{f}$ contains the point $f_{L}(z)$ provided that $f_{L}$ is defined on $z$. It follows that $f_{L}(z)\in J^{L}_{*}$. Note, finally, that $f_{L}$ can be undefined at $z$ only if $z$ is one of the finitely many vertices of $\Gamma_{1,L}$. Finally, the fact that points of $J^{L}_{*}$ are nonwandering under the iterates of $f_{L}$ follows from the corresponding property of $h_{f}$. ∎
@@ 6.2. Geometry and dynamics of $J^{L}_{*}$
In this section, we discuss what the set $J^{L}_{*}$ may look like and how $f_{L}$ may act on this set. Assume that $J^{L}_{*}\neq\varnothing$ for a given choice of a co-oriented line $L$. We think of $L$ as being a horizontal line, and will refer to the positive side of $L$ as the upper side. Consider the union $I_{L}$ of all edges of $\Gamma_{1,L}$ that contain points of $J^{L}_{*}$ inside, i.e., not only as endpoints. Then $f_{L}$ acts on $I_{L}$ as a translation scissor congruence and takes $I_{L}$ into $\Gamma_{1,L}$. It can be extended to an interval exchange transformation $\varphi_{L}$ of $\Gamma_{1,L}$. Note that the dynamics of $\varphi_{L}$ and $f_{L}$ must match on $J^{L}_{*}$ but may be different otherwise. We need some general properties of interval exchange maps.
@@
Let $\varepsilon>0$ be a sufficiently small real number, in particular, smaller than the distance between any two distinct lines containing components of $J^{L}_{*}$. A forward trapezoid is defined as the trapezoid with horizontal bases such that the bottom base is a forward 1-cell of $J^{L}_{*}$, the height is $\varepsilon$, and, near the endpoints of its bottom base, the trapezoid coincides with elementary sectors. When $\varepsilon$ needs to be indicated, we speak of forward $\varepsilon$-trapezoids. Backward trapezoids are defined in the same way, replacing forward 1-cells with backward 1-cells. The union of the forward trapezoids is called the forward $\varepsilon$-train of $J_{L*}$, or just the forward train of $J_{L*}$ if $\varepsilon$ is fixed. Backward trains are defined similarly.
@@
Note that the $\varepsilon$-trains of $J^{L}_{*}$, both forward and backward, are contained in an arbitrarily small neighborhood of $J^{L}_{*}$, provided that $\varepsilon$ is small enough for the chosen neighborhood. The set theoretic difference between the backward train and the forward train, with the same $\varepsilon$, consists of several triangles (possibly degenerate, i.e., line segments) cut out of the union of the two trains by $\Gamma_{1}$. Similarly, the forward train minus the backward train is a union of finitely many triangles cut out of the union of the two trains by $\Gamma_{-1}$.
@@
If $f$ is unbranched, then there can be no non-degenerate triangles cut out by $\Gamma_{1}$ or by $\Gamma_{-1}$. Therefore, the forward and the backward trains are subdivisions of the same polygon, and $f$ acts on this polygon as a polygon exchange transformation. Moreover, the restriction of this polygon exchange transformation to every component of its domain is a parallel translation by a vertor parallel to $L$, that is, by a horizontal vector.
@@
Suppose that an essential nest $(Q_{n})$ belongs to $\mathcal{J}_{*}$, and fix $\varepsilon>0$. There is a co-oriented line $L$ parallel to an edge of $\Gamma_{1}$ and such that, for all sufficiently large $n$, the bi-piece $Q_{n}$ is contained in the $\varepsilon$-train of $J^{L}_{*}$.
@@ 6.3. Horizontal polygon exchange transformations
Fix a co-oriented line $L$ and, as before, refer to its direction as the horizontal direction, and to its positive side as the upper side. Suppose that a polygon exchange transformation $f_{\Theta}$ of a polygon $\Theta$ translates all level 1 pieces by vectors parallel to $L$. In this case, $f_{\Theta}$ is called a horizontal polygon exchange transformation with respect to $L$. Our standing assumption in Section 6.3 is that $f_{\Theta}$ is a horizontal polygon exchange transformation with respect to $L$. There are several special properties of $f_{\Theta}$. Any horizontal line $L_{t}$ parallel to $L$ intersects $\Theta$ over a finite union of horizontal intervals. Note that $f_{\Theta}$ induces an interval exchange transformation of $\Theta\cap L_{t}$. Here, the parameter $t$ can be understood, e.g., as a height over $L=L_{0}$, and we write $\varphi_{t}$ as the interval exchange transformation of $\Theta\cap L_{t}$ induced by $f_{\Theta}$.
@@
Before giving a proof of Theorem 6.13, we make several preparatory remarks. There are finitely many lines $L_{t_{0}}$, $\dots$, $L_{t_{m}}$ such that the vertices of all level 1 pieces of $f_{\Theta}$ are contained in the union of these lines. Assume that the set of lines $L_{t_{i}}$ with the just mentioned property is chosen to be minimal by inclusion. In this case, the part $\Theta_{i}$ of $\Theta$ between $L_{t_{i}}$ and $L_{t_{i+1}}$ is a union of trapezoids with horizontal edges in $L_{t_{i}}$ and $L_{t_{i+1}}$. Some of these edges may degenerate to points, letting the corresponding trepezoids degenerate to triangles. Call the thus obtained trapezoids between $L_{t_{i}}$ and $L_{t_{i+1}}$ the trapezoids on the $i$-th floor. Observe that $f_{\Theta}$ acts as a horizontal polygon exchange transformation on $\Theta_{i}$. Knowing that the dynamic invariant $\mathrm{Inv}_{L}$ is additive with respect to polygonal subdivisions of the domain, it suffices to prove Theorem 6.13 for the union of all trapezoids on a given floor. Moreover, if we consider the first return map to any given trapezoid and then subdivide it further into floors if necessary, then we reduce Theorem 6.13 to the case when $\Theta$ is just an $L$-trapezoid, i.e., a trapezoid with top and bottom edges, possibly degenerate, parallel to $L$. We therefore assume now that $\Theta$ is an $L$-trapezoid.
@@
where $x\wedge y$ means $x\otimes y-y\otimes x$, and all tensor products are over $\mathbb{Q}$. Here, we used that $\lambda_{i}\otimes\Omega_{ij}\lambda_{j}=\Omega_{ij}\lambda_{i}\otimes\lambda_{j}$, since $\Omega_{ij}\in\mathbb{Q}$, and $\otimes$ is bilinear over $\mathbb{Q}$.
@@
Write $\lambda_{i}(t)$ as $\lambda_{i}(0)+t\lambda^{\prime}_{i}(0)$ and substitute this expression into the formula $\sum\lambda_{i}(t)\otimes\tau_{i}$ for the invariant $\mathrm{Inv}(\varphi_{t})$. Recall that $\mathbb{R}\otimes\mathbb{R}$ can be viewed as a vector space over $\mathbb{R}$, with left multiplication by real scalars satisfying the identity $(ta)\otimes b=t(a\times b)$. It follows that $\mathrm{Inv}(\varphi_{t})$ can be written as the sum of two terms: the first term is independent of $t$ and is equal to $\mathrm{Inv}(\varphi_{0})$, whereas the second term is $t$ times the sum of $\lambda^{\prime}_{i}(0)\otimes\tau_{i}$ over $i$ ranging from 1 to $m$. Clearly, it suffices to show that this second term, equal to $\frac{d}{dt}\mathrm{Inv}(\varphi_{t})|_{t=0}$, vanishes (the derivative makes sense since $\varphi_{t}$ takes values in a two-dimensional $\mathbb{R}$-vector subspace of $\mathbb{R}\otimes\mathbb{R}$). Substituting the expression for $\tau$ as $\tau=\Omega\lambda(0)$, we obtain that
@@
Here, the factor $\Omega_{ij}$ has been moved to the front of the tensor product since $\Omega_{ij}\in\mathbb{Q}$.
@@
By the above, it suffices to assume that $\Theta$ is an $L$-trapezoid between $L_{0}$ and $L_{1}$, so that $\varphi_{t}$ for $t\in[0;1]$ form a linear one parameter family of interval exchange transformations. Since all the translation vectors are horizontal, the space $\mathbb{R}^{2}$ of translations reduces to $\mathbb{R}^{1}$, where all $\tau_{i}(t)$ lie. With this identification, the dynamic Hadwiger invariant of $f_{\Theta}$ identifies with the difference $\mathrm{Inv}(\varphi_{0})-\mathrm{Inv}(\varphi_{1})$ of two SAF invariants. By Lemma 6.15, the latter difference is zero. ∎
As it turns out, other dynamic Hadwiger invariants of $f_{\Theta}$ vanish as well.
@@
All dynamic Hadwiger invariants of a horizontal polygon exchange transformation $f_{\Theta}$ are zero.
@@
Vanishing of $\mathrm{Inv}_{L}(f_{\Theta})$ follows from Theorem 6.13. Let $L^{\prime}$ be a line transverse to $L$; we want to show that $\mathrm{Inv}_{L^{\prime}}(f_{\Theta})=0$. It suffices to assume that $\Theta$ is an $L$-trapezoid between $L_{0}$ and $L_{1}$. Also, one may assume without loss of generality that the positive side of $L^{\prime}$ is on the right of $L^{\prime}$. Write $\ell^{\prime}$ for the length of the intersection of the strip between $L_{0}$ and $L_{1}$ with any line parallel to $L^{\prime}$. Suppose that $i_{1}<\dots<i_{k}$ are all the values of the index $i$, for which the left end of $I_{i}(t)$ for all $t\in(0;1)$ lies on an edge of $\Gamma_{1}(f_{\Theta})$ parallel to $L^{\prime}$. One also writes $j_{1}<\dots<j_{l}$ for the values of $j$ such that the right end of $I_{j}(t)$ for all $t\in(0;1)$ lies on an edge of $\Gamma_{1}(f_{\Theta})$ parallel to $L^{\prime}$. The value $\mathrm{Inv}_{L^{\prime}}(f_{\Theta})$ is by definition equal to
@@
All the dynamic Hadwiger invariants of $f_{\circ}$ are the same as those of $f$. Hence, at least one of these invariants takes a nonzero value on $f_{\circ}$.
@@
Indeed, this follows from the additivity of the invariants and the fact that the restriction of $f$ to every train has all the dynamic Hadwiger invariants zero, by Theorem 6.16. ∎
Since $f_{\circ}$ has a nonzero dynamic Hadwiger invariant, it is not periodic. However, $f_{\circ}$ cannot have aperiodic points — all aperiodic points of $f_{\circ}$ would also be aperiodic points of $f$. The ideal Julia set $\mathcal{J}(f_{\circ})$ of $f_{\circ}$ and the essential part of it, $\mathcal{J}_{*}(f_{\circ})$, are therefore nonempty.